# OccFiner: Offboard Occupancy Refinement with Hybrid Propagation

### v1

@@ 3.3 Region-centric Global Propagation
Vanilla Registration: As illustrated in Fig. 4, it aggregates independent SSC predictions from $12n+$ 1 frames into a multi-view consistent prediction for each region, where $n$ is the temporal window radius. This involves the aggregation of per-frame semantics and geometry $\{\hat{Y}_{i}\}_{i=1}^{2n+1}$ into a refined SSC map. For a given target voxel index $\mathbf{y}$ within a 3D region of target $\hat{Y}_{tg}$, the process starts with devoxelizing the current SSC voxel $\hat{Y}_{cur}$ to derive the semantic point cloud $\mathbf{P}_{cur}$ in the current LiDAR coordinate system:
$\mathbf{P}_{cur}(x,y,z,c)=\left\{\left(o_{x}+i\cdot\text{dv}_{x},o_{y}+j\cdot\text{dv}_{y},o_{z}+k\cdot\text{dv}_{z},c\right)\,|\,\hat{Y}_{ijk}=c\right\},$ (7)
where the voxel origin is $(o_{x},o_{y},o_{z})$. dv denotes the voxel size and $c$ represents the semantic classes. This point cloud is then transformed relative to the target LiDAR coordinate system with poses $\{{T}_{cur}^{tg}\}_{tg-n}^{tg+n}$:
@@
Incorporating with Sensor Bias: While vanilla registration averages SSC across multiple frames, it overlooks the differential weighting of different locations. To address this, sensor-aware weighting is introduced, tailored to the distinct characteristics of cameras and LiDARs. This approach is informed by the measurement characteristics of measurement devices like cameras and LiDARs. Camera systems are constrained by their field of view, with points within the viewing frustum yielding higher accuracy compared to those outside. Additionally, the occlusion of distant points by nearer ones necessitates lower voting weights for the former. Conversely, LiDAR systems exhibit increased sparsity and depth measurement errors at longer distances. Therefore, for cameras, points within the field of view and at closer ranges are given higher voting weights:
$F_{cam}(x_{c},y_{c},z_{c})=\left(|\theta_{xc}|\leq\frac{fov_{h}}{2}\right)\land\left(|\theta_{yc}|\leq\frac{fov_{w}}{2}\right)\land(z_{c}>0),$ (10)
$W_{cam}(\mathbf{x})=\begin{cases}w_{\text{high}},&\text{if $(F_{cam}\land B)(\mathbf{x})>0$}\\ w_{\text{med}},&\text{if $[F_{cam}(\mathbf{x})>0]$ $\land$ $[B(\mathbf{x})<0]$}\\ w_{\text{low}},&\text{otherwise}\end{cases}$ (11)
where $(x_{c},y_{c},z_{c})$ are the voxel indices in the camera coordinate system, $fov$ denotes the field of view of the camera, $F_{cam}$ represents the voxels within the frustum, and $B$ is the bounding box voxel with values being True inside the box and False outside. $\mathbf{x}$ is the vertex and $W_{cam}$ is the final camera voting weight.
While for LiDARs, the voting weight linearly attenuates with increasing distance from the LiDAR center:
$W_{li}(r)=w_{\text{max}}-(w_{\text{max}}-w_{\text{min}})\times\frac{r}{R},$ (12)
where $r$ denotes the radial distance from the origin in the LiDAR coordinate system, $R$ represents the maximum range, and $w_{\text{max}}$ and $w_{\text{min}}$ correspond to the maximum and minimum weights, respectively. Consequently, the OccFiner framework adopts a hybrid dual-stage strategy: the first stage addresses onboard model errors, enhancing prediction accuracy regardless of the range, and the second stage adjusts for the physical limitations of cameras and LiDARs, further improving the offboard accuracy.
3.4 Supervision
Our multi-to-multi local propagation network is trained with the labels from the training set frame-by-frame, while the region-centric global propagation is a learning-free process. The SSC loss $\mathcal{L}_{ssc}$ for network training consists of two parts as below:
$\mathcal{L}_{ssc}=\mathcal{L}_{ce}+\mathcal{L}_{lovasz},$ (13)
where $\mathcal{L}_{ce}$ is the cross-entropy loss and $\mathcal{L}_{lovasz}$ is the Lovász loss [3].
4 Experiments
4.1 Datasets and Implementation Details
SemanticKITTI and SSCBench-KITTI360: The SemanticKITTI dataset [2] is a pivotal SSC benchmark featuring $22$ outdoor driving scenarios, segmented into train, validation, and test sets with a $10/1/11$ split. The focus is on a specific volume around the vehicle, mapped in $256\times 256\times 32$ voxel grids, where each voxel measures $0.2m^{3}$. SSCBench-KITTI360 [26], derived from KITTI-360 [30], spans $73.7km$ with extensive image and laser scan data, providing diverse geographical coverage and extended sequences. This dataset, sampled to reduce redundancy, offers around $13k$ frames, serving as a vital resource for SSC research.
4.2 Comparison with State-of-the-Art
@@
Table 1: Camera-based semantic scene completion results on the SemanticKITTI validation set [2]. * represents these methods are adapted for the RGB inputs, which are implemented and reported in MonoScene [6]. normal-$\dagger$ represents the reproduced result from [18].
@@
As our work presents the first offboard SSC generation, there are no existing competitors to parallel OccFiner’s performance. Additionally, our OccFiner can utilize any off-the-shelf onboard model as its intermediate component. As shown in Tab. 1, we use the camera-based SSC algorithm [6, 18, 27] as the onboard component. On the SemanticKITTI validation set [2], OccFiner demonstrates substantial improvements in mIoU accuracy for various algorithms like MonoScene [6], TPVFormer [18], and VoxFormer [27]. The mIoU accuracy is relatively improved by $21.39\%$, $17.08\%$, and $35.71\%$, respectively. Notably, OccFiner (Vox) achieves the new state-of-the-art camera SSC accuracy of $18.09\%$ mIoU. To ensure a meaningful and fair comparison, we also establish an offboard baseline algorithm, “Average Fusion”. It performs region-centric aggregation without learning refinement or considering sensor measurement biases. Despite this, compared to VoxFormer + Average Fusion, our OccFiner (Vox) can still improve SSC quality by a large margin of over $24.59\%$ in mIoU under the long-range setting.
Qualitative comparisons. As shown in Fig. 5, our offboard solution, OccFiner, effectively corrects major onboard errors, notably in resolving extensive missing road and vehicle voxels. Furthermore, OccFiner enhances the effectiveness of purely visual approaches, surpassing the traditional LiDAR-based onboard method SSCNet-full [55] in performance, indicating its potential to advance the field of autonomous navigation with pure visual solution.
Table 2: Quantitative comparison against the state-of-the-art LiDAR-based SSC methods on the SemanticKITTI validation set. normal-$\dagger$ results from the official code release.
Methods Modality IoU (%) mIoU (%) Rel.
12.8m 25.6m 51.2m 12.8m 25.6m 51.2m
@@
Our exploration extends to validating OccFiner’s versatility across different modalities in offboard SSC generation. As shown in Tab. 2, for LiDAR-based SSC models like SSCNet [55], SSCNet-full [55], LMSCNet [52], and SCPNet [65], OccFiner consistently attains relative mIoU improvements of $36.64\%$, $20.92\%$, $15.65\%$, and $14.83\%$, respectively. This demonstrates its modality-agnostic capability. We observe that the enhancements offered by OccFiner are consistent across varying distances. This improvement is not only confined to long distances and low-accuracy scenarios but also extends to challenging high-accuracy short-range as well, indicating a robust and versatile perception of the framework. Remarkably, the accuracy of OccFiner (VoxFormer) even surpasses onboard LiDAR-based SSC models including SSCNet, SSCNet-full, and LMSCNet. This is the first time that the camera-based SSC model surpasses the LiDAR-based SSC model in accuracy. Compared with the onboard LiDAR-based LMSCNet, the camera-based OccFiner (VoxFormer) achieves a relative mIoU improvement of $5.24\%$. This proves that OccFiner unleashes the potential of the camera-based SSC applications. This breakthrough in the camera-based versus LiDAR-based SSC model signifies a crucial advancement for autonomous driving, notably in the realms of pure visual SSC data closure and auto-labeling.
Table 3: Semantic scene completion results on the SemanticKITTI hidden test set [2]. * represents these methods are adapted for the RGB inputs, which are implemented and reported in MonoScene [6]. normal-$\dagger$ results from the official code release.
@@
As shown in Tab. 3, in the results submitted for the hidden test set of SemanticKITTI [2], OccFiner showcases remarkable performance. The OccFiner (Vox) and OccFiner (SCPNet) establish new state-of-the-art accuracy for camera-based and LiDAR-based SSC respectively. They deliver significant relative mIoU improvements compared to the best-published results of $26.99\%$ and $9.50\%$, respectively. The OccFiner (Vox) consistently surpasses the SSCNet-full [55]. This further demonstrates the efficacy and advancement of the OccFiner framework.
Table 4: Benchmarking results on SSCBench-KITTI360 [26]. The default evaluation range is $51.2{\times}51.2{\times}6.4m^{3}$.
@@
The effectiveness of OccFiner is also assessed on the SSCBench-KITTI360 dataset [26]. As shown in Tab. 4, OccFiner (Mono) demonstrates notable performance, achieving a relative improvement in mIoU of $7.96\%$ compared to MonoScene [6]. This underscores OccFiner’s capability to enhance SSC accuracy across diverse datasets.
4.3 Ablation Studies

### v4

@@ III-C Region-centric Global Propagation
Vanilla Registration: As illustrated in Fig. 4, it aggregates independent SSC predictions from $12n+$ 1 frames into a multi-view consistent prediction for each region, where $n$ is the temporal window radius. This involves the aggregation of per-frame semantics and geometry $\{\hat{Y}_{i}\}_{i=1}^{2n+1}$ into a refined SSC map. For a given target voxel index $\mathbf{y}$ within a 3D region of target $\hat{Y}_{tg}$, the process starts with devoxelizing the current SSC voxel $\hat{Y}_{cur}$. Let $(o_{x},o_{y},o_{z})$ be the coordinate origin of the voxel grid, $\text{dv}=(\text{dv}_{x},\text{dv}_{y},\text{dv}_{z})$ be the size of each voxel, and $c$ be a semantic class label. For each occupied voxel $\hat{Y}_{ijk}=c$ at index $(i,j,k)$, we derive a point in the semantic point cloud $\mathbf{P}_{cur}$ in the current LiDAR coordinate system as follows:
$\mathbf{P}_{cur}(x,y,z,c)$ =$=\left\{\left(o_{x}+i\cdot\text{dv}_{x},o_{y}+j\cdot\text{dv}_{y},\right.\right.$ (8)
$\left.\left.o_{z}+k\cdot\text{dv}_{z},c\right)\,|\,\hat{Y}_{ijk}=c\right\}.$
Next, this point cloud is then transformed relative to the target LiDAR coordinate system with poses $\{{T}_{cur}^{tg}\}_{tg-n}^{tg+n}$:
@@
Incorporating with Sensor Bias: While vanilla registration averages SSC across multiple frames, it overlooks the differential weighting of different locations. To address this, sensor-aware weighting is introduced, tailored to the distinct characteristics of cameras and LiDARs. This approach is informed by the measurement characteristics of measurement devices like cameras and LiDARs. Camera systems are constrained by their field of view, with points within the viewing frustum yielding higher accuracy compared to those outside. Additionally, the occlusion of distant points by nearer ones necessitates lower voting weights for the former. Conversely, LiDAR systems exhibit increased sparsity and depth measurement errors at longer distances.
Therefore, for cameras, points within the field of view ($fov$) and closer ranges are given higher voting weights. Let $F_{cam}(x_{c},y_{c},z_{c})$ be a boolean function indicating if a voxel is within the camera frustum, defined as:
[…]
$\land(z_{c}>0),$
where $(x_{c},y_{c},z_{c})$ denote the voxel indices in the camera coordinate system. Let $B(\mathbf{x})$ be a boolean function indicating if a voxel vertex $\mathbf{x}$ is inside a predefined bounding box (used to prioritize closer regions). We define three weight levels: high ($w_{\text{high}}$), medium ($w_{\text{med}}$), and low ($w_{\text{low}}$). The final camera voting weight $W_{cam}(\mathbf{x})$ for a voxel vertex $\mathbf{x}$ is assigned as:
$W_{cam}(\mathbf{x})=\begin{cases}w_{\text{high}},&\text{if $(F_{cam}\land B)(\mathbf{x})>0$};\\ w_{\text{med}},&\text{if $[F_{cam}(\mathbf{x})>0]$ $\land$ $[B(\mathbf{x})<0]$};\\ w_{\text{low}},&\text{otherwise}.\end{cases}$ (12)
Conversely, for LiDARs, accuracy typically decreases with distance due to sparsity and measurement errors. Let $r$ be the radial distance of a voxel from the LiDAR origin, $R$ be the maximum considered range, and $w_{\text{max}}$ and $w_{\text{min}}$ be the maximum and minimum voting weights. The voting weight $W_{li}(r)$ linearly attenuates with increasing distance $r$:
$W_{li}(r)=w_{\text{max}}-(w_{\text{max}}-w_{\text{min}})\times\frac{r}{R}.$ (13)
@@
TABLE II: Camera-based semantic scene completion results on the SemanticKITTI validation set [16]. * represents these methods are adapted for the RGB inputs, which are implemented and reported in MonoScene [1]. $\dagger$ represents the reproduced result from [81].
@@
Consequently, the OccFiner framework adopts a hybrid dual-stage strategy: the first stage addresses onboard model errors, enhancing prediction accuracy regardless of the range, and the second stage adjusts for the physical limitations of cameras and LiDARs, further improving the offboard accuracy.
III-D Supervision
The multi-to-multi local propagation network is trained end-to-end using the ground-truth labels corresponding to the input frames. The overall loss function for Semantic Scene Completion ($\mathcal{L}_{ssc}$) comprises two commonly used components. The first is the standard voxel-wise Cross-Entropy loss, denoted as ($\mathcal{L}_{ce}$). The second is the Lovász loss [84], denoted as $\mathcal{L}_{lovasz}$, which is designed as a surrogate for directly optimizing the intersection-over-union (IoU) metric crucial for segmentation tasks. The total SSC loss is the sum of these two components:
$\mathcal{L}_{ssc}=\mathcal{L}_{ce}+\mathcal{L}_{lovasz}.$ (14)
IV Experiments
TABLE III: Quantitative comparison against the state-of-the-art LiDAR-based SSC methods on the SemanticKITTI validation set. $\dagger$ results from the official code release. The ‘Relative Improvement’ column shows the relative percentage improvement in the overall $51.2m\times 51.2m\times 6.4m$ mIoU achieved by OccFiner compared to its corresponding onboard baseline model listed directly above it.
Methods Modality IoU (%) mIoU (%) Relative
12.8m 25.6m 51.2m 12.8m 25.6m 51.2m Improvement
@@ IV Experiments
IV-A Datasets and Implementation Details
SemanticKITTI and SSCBench-KITTI360: The SemanticKITTI dataset [16] is a pivotal SSC benchmark featuring $22$ outdoor driving scenarios, segmented into train, validation, and test sets with a $10/1/11$ split. The focus is on a specific volume around the vehicle, mapped in $256\times 256\times 32$ voxel grids, where each voxel measures $0.2m^{3}$. SSCBench-KITTI360 [68], derived from KITTI-360 [85], spans $73.7km$ with extensive image and laser scan data, providing diverse geographical coverage and extended sequences. This dataset, sampled to reduce redundancy, offers around $13k$ frames, serving as a vital resource for SSC research. Unless otherwise specified (e.g., distance-based breakdown in Tab. III), all reported IoU and mIoU metrics are calculated over the standard full evaluation volume of $256\times 256\times 32$ voxels, corresponding to a spatial range of $51.2m\times 51.2m\times 6.4m$ centered around the ego-vehicle.
Details for Multi-to-Multi Local Propagation Network: In our implementation, the Bird’s-Eye View (BEV) feature and positional encoders are structured as an embedding Linear layer with LayerNorm and four convolutional blocks. Each block uses MaxPool for downsampling by a factor of $2$, includes two BEV $3\times 3$ 2D convolutions with ReLU activation, and outputs features with a channel of $80$. For the DualFlow4D block’s BEV branch, we set the bev patch size at $W_{h}=7$, $W_{w}=7$, and $T=6$. The local and distant reference time windows are $T_{l}=4$ and $T_{r}=2$, respectively, with a reference frame sampling temporal radius of $n=10$ and a hidden dimension of $c_{e}=256$. The pillar branch of the DualFlow4D block is set at $W_{z}=32$ and $T=6$. Across all experiments, we stack $N=2$ DualFlow4D Transformer blocks to achieve the desired functionality.
Details for Region-centric Global Propagation: For global propagation in each SSC frame, we use a temporal radius of $n=25$. Camera voting weights are set at $w_{high}=1$, $w_{med}=0.1$, and $w_{low}=0.01$, with a bounding box of $25.6m\times 25.6m\times 6.4m$. For LiDAR, the voting weights range from $w_{max}=10$ to $w_{min}=0.1$. We employ vectorization to expedite the global propagation and, due to memory constraints, use single-precision (float) calculations for semantic voxel registration and 8-bit quantization (uint8) in generating city-level global street view maps. In addition, when determining the semantic category of each voxel of the city-level SSC map, we use a three-dimensional sliding window strategy to implement argmax for different map chunks and finally fuse them into a global three-dimensional semantic map to further reduce memory consumption.
TABLE IV: Semantic scene completion results on the SemanticKITTI hidden test set [16]. * represents these methods are adapted for the RGB inputs, which are implemented and reported in MonoScene [1]. $\dagger$ results from the official code release.
@@ IV-A Datasets and Implementation Details
TABLE V: Benchmarking results on SSCBench-KITTI360 [68]. The default evaluation range is $51.2{\times}51.2{\times}6.4m^{3}$.
@@
IV-B Comparison with State-of-the-Art Methods
As our work presents the first offboard SSC generation, there are no existing competitors to parallel OccFiner’s performance. Additionally, our OccFiner can utilize any off-the-shelf onboard model as its intermediate component.
Quantitatively Comparisons: As shown in Tab. II, we use the camera-based SSC algorithm [1, 81, 4] as the onboard component. On the SemanticKITTI validation set [16], OccFiner demonstrates substantial improvements in mIoU accuracy for various algorithms like MonoScene [1], TPVFormer [81], and VoxFormer [4]. The mIoU accuracy is relatively improved by $21.39\%$, $17.08\%$, and $35.71\%$, respectively. Notably, OccFiner (Vox) achieves the new state-of-the-art camera SSC accuracy of $18.09\%$ mIoU. To ensure a meaningful and fair comparison, we also establish an offboard baseline algorithm, “Average Fusion”. It performs region-centric aggregation without learning refinement or considering sensor measurement biases. Despite this, compared to VoxFormer + Average Fusion, our OccFiner (Vox) can still improve SSC quality by a large margin of over $24.59\%$ in mIoU under the long-range setting.
TABLE VI: Auto-labeling comparison between onboard models trained with either training set ground-truth labels (GT-Train) or auto-labeling mixed GT-Train + pseudo-labels (PL) generated by different sources.
Label Source IoU mIoU
GT-Train Human Labeling 36.42 13.17
GT-Train + PL-Test (SSCNet [18]) LiDAR Auto-Labeling 37.07 13.63
GT-Train + PL-Test (VoxFormer [4]) Visual Auto-Labeling 36.15 12.74
GT-Train + PL-Test (VoxFormer [4] + OccFiner) Visual Auto-Labeling 37.17 13.90
TABLE VII: City-level semantic scene completion results on the SemanticKITTI validation set [16]. Note we only compare static classes in the global map comparisons. $\dagger$ results from the official code release.
Method Modality
$\blacksquare$ road (15.30%)
$\blacksquare$ parking (1.12%)
$\blacksquare$ sidewalk (11.13%)
$\blacksquare$ other-ground (0.56%)
$\blacksquare$ building (14.1%)
$\blacksquare$ fence (3.90%)
$\blacksquare$ vegetation (39.3%)
$\blacksquare$ trunk (0.51%)
$\blacksquare$ terrain (9.17%)
$\blacksquare$ pole (0.29%)
$\blacksquare$ traffic-sign (0.08%)
IoU mIoU
SSCNet [18] LiDAR 32.78 11.88 18.00 0.00 14.58 9.17 26.78 13.48 26.56 19.88 4.70 30.06 16.16
SSCNet-full [18] LiDAR 43.82 12.52 21.24 0.35 14.58 10.05 27.18 16.46 33.25 24.47 6.92 32.83 19.16
LMSCNet [2] LiDAR 49.71 16.73 25.89 0.00 18.93 11.67 34.73 15.68 35.16 26.81 0.69 41.66 21.45
[…]
OccFiner (SCP) LiDAR 72.76 51.56 50.57 10.03 41.40 30.61 47.09 39.86 55.11 42.64 26.76 60.16 42.58
w.r.t. SCPNet - -2.05 -0.87 -0.23 +2.18 +2.49 +1.62 +1.79 +3.59 +0.25 +1.31 +1.17 +0.77 +1.03
MonoScene [1] Camera 47.07 10.04 15.32 0.58 10.39 9.74 17.41 5.13 25.44 7.31 5.37 29.75 13.98
OccFormer [3] Camera 46.70 17.90 14.77 0.95 9.17 7.31 20.40 5.93 24.64 6.87 5.77 27.63 14.58
TPVFormer [81] Camera 43.49 15.28 13.46 0.51 8.91 8.74 17.28 4.86 24.53 6.34 4.23 26.61 13.42
VoxFormer [4] Camera 42.22 12.48 15.16 0.36 13.04 10.12 22.21 10.02 25.63 12.36 6.71 32.66 15.48
OccFiner (Vox) Camera 47.44 16.22 21.04 0.83 13.29 9.96 27.17 10.62 24.75 10.33 5.19 33.59 16.99
w.r.t. VoxFormer - +5.22 +3.74 +5.88 +0.47 +0.25 -0.16 +4.96 +0.60 -0.88 -2.03 -1.52 +0.93 +1.51
Our exploration extends to validating OccFiner’s versatility across different modalities in offboard SSC generation. As shown in Tab. III, for LiDAR-based SSC models like SSCNet [18], SSCNet-full [18], LMSCNet [2], and SCPNet [17], OccFiner consistently attains relative mIoU improvements of $36.64\%$, $20.92\%$, $15.65\%$, and $14.83\%$, respectively. This demonstrates its modality-agnostic capability. We observe that the enhancements offered by OccFiner are consistent across varying distances. This improvement is not only confined to long distances and low-accuracy scenarios but also extends to challenging high-accuracy short-range as well, indicating a robust and versatile perception of the framework. Remarkably, the accuracy of OccFiner (VoxFormer) surpasses several widely recognized onboard LiDAR-based SSC models, including SSCNet [18], SSCNet-full [18], and LMSCNet [2] (comparing Tab. II results with Tab. III). This marks a milestone where a refined camera-based occupancy system outperforms these established LiDAR-based approaches in overall mIoU. Compared with the onboard LiDAR-based LMSCNet, the camera-based OccFiner (VoxFormer) achieves a relative mIoU improvement of $5.24\%$. This proves that OccFiner unleashes the potential of the camera-based SSC applications. This breakthrough in the camera-based versus LiDAR-based SSC model signifies a crucial advancement for autonomous driving, notably in the realms of pure visual SSC data closure and auto-labeling.
As shown in Tab. IV, in the results submitted for the hidden test set of SemanticKITTI [16], OccFiner showcases remarkable performance. The OccFiner (Vox) and OccFiner (SCPNet) establish new state-of-the-art accuracy for camera-based and LiDAR-based SSC respectively. They deliver significant relative mIoU improvements compared to the best-published results of $26.99\%$ and $9.50\%$, respectively. The OccFiner (Vox) consistently surpasses the SSCNet-full [18]. This further demonstrates the efficacy and advancement of the OccFiner framework.
@@ IV-B Comparison with State-of-the-Art Methods
The effectiveness of OccFiner is also assessed on the SSCBench-KITTI360 dataset [68]. As shown in Tab. V, OccFiner (Mono) demonstrates notable performance, achieving a relative improvement in mIoU of $7.96\%$ compared to MonoScene [1]. This underscores OccFiner’s capability to enhance SSC accuracy across diverse datasets.
Qualitative Comparisons: As shown in Fig. 5, our offboard solution, OccFiner, effectively corrects major onboard errors, notably in resolving extensive missing road and vehicle voxels. Furthermore, OccFiner enhances the effectiveness of purely visual approaches, surpassing the traditional LiDAR-based onboard method SSCNet-full [18] in performance, indicating its potential to advance the field of autonomous navigation with pure visual solution.
Latency: The online component VoxFormer [4] exhibits a latency of $1.52$ s/frame on a single NVIDIA RTX 3090 with single-precision computation. In contrast, the multi-to-multi local propagation network of OccFiner incurs a computational delay of $0.027$ s/frame on the same device. Furthermore, the region-centric global propagation on an Intel i9-13980HX CPU has a computational delay of 5.96 s/frame. The latency of OccFiner’s Stage 1 network primarily depends on its own architecture and the input voxel grid size. The Stage 2 latency depends on implementation, and the number of frames being aggregated (temporal window size, $n$). These timings for the VoxFormer baseline illustrate the typical computational overhead introduced by OccFiner’s distinct offboard stages: a minimal increase from Stage 1 suited for near-real-time refinement if needed, and a larger latency for the comprehensive offboard aggregation in Stage 2.
Analysis of Per-Class Performance: As shown in Tab. II, OccFiner (Vox) improves overall IoU/mIoU but shows slight declines for small objects (motorcycle, poles) versus VoxFormer [4]. This stems from offboard multi-view temporal aggregation and temporal smoothing: although highly effective for enhancing the consistency and completeness of larger structures by leveraging consensus across frames, this aggregation may occasionally average out or fail to reinforce the sparse and potentially inconsistent predictions associated with small, thin, or infrequently observed objects, especially if the onboard model’s predictions for these classes are already weak. The trade-off between scene coherence and rare-class fidelity warrants future study.
IV-C Auto-labeling Experiments
Figure 6: City-level semantic scene completion. The city-scale scene is expanded from the original size of $256\times 256\times 32$ to $2205\times 4296\times 261$.
Figure 7: City-level semantic scene completion in BEV. Only static classes are considered. Our proposed OccFiner preserves more high-quality details and reduces camera frustum artifacts.
The aforementioned results highlight OccFiner’s ability to generate superior Semantic Scene Completion (SSC) labels, indicating its promise as an offboard solution for automatic SSC labeling and data closure, particularly for augmenting manual labeling in novel scenarios using pure vision-based solution. To evaluate OccFiner’s label quality further, we conduct auto-labeling experiments on SemanticKITTI [16], detailed in Table VI, which simulate a real-world data loop. Utilizing manually annotated labels as a base, we enrich new scenes from the test sequences with pseudo-labels created by the pure-vision VoxFormer + OccFiner. Comparing training results of an onboard visual SSC model [3] from scratch under different labeling conditions, we find that directly using VoxFormer for ‘Visual Auto-Labeling’ cannot achieve data closure, and the IoU and mIoU of the onboard model dropped by $0.27\%$ and $0.43\%$, respectively. Meanwhile, ‘LiDAR Auto-Labeling’ based on SSCNet [18] can improve the accuracy of the visual onboard model by $0.46\%$. Keeping hyperparameters and model structure consistent, we find that OccFiner’s ‘Visual Auto-Labeling’ outperforms ‘Human Labeling’ alone. Consequently, the onboard model’s IoU on the validation set improved by $0.75$%, with a corresponding large increase of $0.73$% in mIoU. When compared under varied labeling conditions, OccFiner’s visual auto-labeling significantly improves onboard model performance, demonstrating a notable advantage over both pure human labeling and even LiDAR-based auto-labeling. This underscores our method’s ability to effectively close data loops in SSC applications using high-quality, visually-derived labels.